(TCO 2) Find the velocity of a plane that travels 0.600 megamiles due north in 3 hours and 15 minutes.
Answer #1

Distance traveled = 0.6 megamiles = 0.6*106 miles = 6*105 miles

Time taken by plane: t = 3 hours 15 minutes = 3.25 hours

Speed = distance/time = 6*105 miles/3.25 hours = 184615.38 miles/hour = 184615.38*.44704 m/s = 82530.46 m/s